# Cost-aware Stopping for Bayesian Optimization

### v2

@@ 1 Introduction
In this work, we study how to design cost-aware stopping rules, motivated by two primary factors. First, state-of-the-art cost-aware acquisition functions such as the Pandora’s Box Gittins Index (PBGI) (Xie et al., 2024) and log expected improvement per cost (LogEIPC) (Ament et al., 2023) have not yet been studied in the adaptive stopping setting. This is important because—as our experiments in Section˜4 will show—for best performance, one should pair different acquisition functions with different stopping rules. Second, while certain stopping rules, such as UCB-LCB (Makarova et al., 2022), are guaranteed to achieve a low simple regret, they are not necessarily guaranteed to do so with low evaluation costs. This is important because—as our experiments will show—UCB-LCB will often incur high evaluation costs, resulting in a high cost-adjusted simple regret.
@@ 3.1 Theoretical Guarantees on Cost-Adjusted Simple Regret
Let $X$ be compact, and let $f:X\to\mathbb{R}$ be a random function with constant mean $\mu$. Consider a Bayesian optimization algorithm that begins at some initial point $x_{1}\in X$ with cost $C=c(x_{1})$, acquires additional points using either the PBGI or LogEIPC acquisition function, and uses the PBGI/LogEIPC stopping rule. Let $\tau=\min_{t\geq 1}\{\sup_{x\in X}\alpha_{t}^{\mathrm{LogEIPC}}(x)\leq 0\}$ be the algorithm’s stopping time, and denote the posterior expected improvement function by $\alpha_{t}^{\mathrm{EI}}(x)=\mathrm{EI}_{f\mid x_{1:t},y_{1:t}}\left(x;y^{*}_{1:t}\right)$. Then, $\alpha_{t}^{\mathrm{EI}}(x_{t})\geq c(x_{t})$ for all $t\leq\tau$.
Proofs are in Appendix˜B. As a consequence of this claim, we show in Theorem˜2 that the PBGI/LogEIPC stopping rule, when paired with PBGI or LogEIPC, achieve strictly better cost-adjusted simple regret than the naive strategy of stopping immediately after the initial evaluation. The last term in (11) is exactly the expected cost-adjusted simple regret of this stopping-immediately baseline ("Immediate"). While this may sound obvious, many acquisition function–stopping rule pairs fail to satisfy this better-than-immediate property in practice (see Figures˜2, 3 and 4), largely because most stopping rules are designed with simple regret alone in mind and do not explicitly account for evaluation costs. To our knowledge, this is the first result establishing a formal guarantee on cost-adjusted simple regret for Bayesian optimization, as opposed to simpler settings like Pandora’s Box.
@@
Consider the setting, acquisition function, and stopping rule specified in Lemma˜1, and let $U=\mu-\mathbb{E}[\min_{x\in X}f(x)]<\infty$. Then the expected cost-adjusted simple regret is bounded by
$\mathbb{E}\left[y_{1:\tau}^{*}-\min_{x\in X}f(x)+\sum_{t=1}^{\tau}c(x_{t})\right]\leq\mu+C-\mathbb{E}\left[\min_{x\in X}f(x)\right]=\mathbb{E}\Big[\underbrace{y_{1}-\min_{x\in X}f(x)+c(x_{1})}_{\textnormal{cost-adjusted regret of "Immediate"}}\Big].$ (11)
@@ 4 Experiments
To evaluate our proposed PBGI/LogEIPC stopping rule, we design three complementary sets of experiments that progressively test its performance. First, we consider an idealized, low-dimensional Bayesian regret setting in which the Gaussian process model exactly matches the true objective. This controlled environment allows us to isolate the effect of different stopping rules without interference from modeling or optimization errors. Then, we move to a higher-dimensional Bayesian regret setting where each acquisition-function minimization must be approximated via a numerical optimizer. Finally, we test in practical scenarios with potential objective model mismatch, using the LCBench hyperparameter tuning benchmark and the NATS neural architecture size search benchmark. In each case, we compare combinations of acquisition functions and stopping rules, which we describe next.
@@
We compare the proposed PBGI/LogEIPC stopping rule against several stopping rules from prior work: UCB-LCB Makarova et al. (2022), LogEIPC-med Ishibashi et al. (2023), SRGap-med Ishibashi et al. (2023), and PRB Wilson (2024). We also include two simple heuristics used in practice: Convergence and GSS. UCB-LCB stops once the gap between upper and lower confidence bounds falls below a configurable threshold $\theta$. LogEIPC-med stops when the log expected improvement per cost drops beneath $\log(\eta)$ plus the median of its initial $I$ values. SRGap-med stops when the gap of the expected minimum simple regret falls below $\chi$ times the median of its initial $I$ values. PRB triggers once a probabilistic regret bound satisfies the regret tolerance $\epsilon$ and confidence $\delta$ parameters. Convergence stops as soon as the best observed value remains unchanged for $w$ iterations. GSS stops if the recent improvement is less than $\phi\times\mathrm{IQR}$ over the past $w$ trials where $\mathrm{IQR}$ denotes the inter-quartile range of current observations. Finally, we include two reference baselines: Immediate, which stops right after the initial evaluation (see Section˜3.1), and Hindsight, which, for each trial, selects the stopping time that yields the lowest cost-adjusted simple regret in hindsight, thus providing a lower bound on achievable performance.
@@ 4.1 Bayesian Regret
Figure 2: Cost-adjusted simple regret across acquisition–stopping rule pairs in 1D Bayesian regret setting. Objectives are sampled from a GP with Matérn-5/2 kernel, with a linear cost function scaled by $\lambda=0.1,0.01,0.001$. Matched PBGI/LogEIPC pairs perform best at $\lambda=0.1,0.01$, and remain near-hindsight-optimal at $\lambda=0.001$, though slightly outperformed by PRB + PBGI/LogEIPC acquisition. The Immediate baseline is omitted at $\lambda=0.001$ due to its significantly higher regret compared to other stopping rules.
Figure 3: Cost-adjusted simple regret across acquisition–stopping rule pairs in 8D Bayesian regret experiments. Objectives are sampled from a GP with a Matérn-5/2 kernel, using three different cost functions scaled by $\lambda=0.01$. The PBGI/LogEIPC stopping rule consistently yields the lowest cost-adjusted regret that is close to hindsight optimal, when paired with LogEIPC, PBGI, and LCB.
@@ 4.2 Empirical Automated Machine Learning Benchmarks
Figure˜4 presents cost-adjusted regret for combinations of acquisition function and stopping rule on classification tasks from LCBench and NATS-Bench, with representative [^3] These were chosen to avoid degenerate cases—neither so large that the policy stops after only a few evaluations (e.g., $\lambda=10^{-4}$ on NATS-Bench, Figure 18) nor so small that evaluations are effectively free. cost-scaling factor $\lambda=10^{-4}$ and $\lambda=10^{-5}$ respectively. Additional results under different cost-scaling factors and the unknown-cost setting are provided in Figures˜15, 16 and 17 of Appendix˜D, along with visual comparisons between adaptive and fixed-iteration stopping. Overall, our PBGI/LogEIPC stopping rule consistently performs strongly in terms of cost-adjusted simple regret, particularly when paired with the PBGI acquisition function. We also report, in Table˜1 of Appendix˜D, how often each stopping rule fails to trigger within the 200-iteration cap: baselines such as SRGap-med and UCB-LCB often fail to stop early—particularly on NATS-Bench—whereas our stopping rule reliably stops before the cap.
@@
Figure 4: Comparison of cost-adjusted simple regret across acquisition function and stopping rule combinations on LCBench and NATS-Bench. The objective is to minimize validation error on classification tasks, with proxy runtime as evaluation cost, scaled by representative values of $\lambda$ ($10^{-4}$ for LCBench and $10^{-5}$ for NATS-Bench); see Figures˜12, 13 and 14 for performance under other $\lambda$. The matched PBGI combination performs best on adult, higgs, and cifar10-valid, second-best on Fashion-MNIST, and third-best on the remaining two, while closely approaching the hindsight optimal across all tasks. Our stopping rule is also competitive when paired with LogEIPC or TS.
@@ Acknowledgments
QX thanks Nairen Cao for helpful discussions on hindsight optimal, Theodore Brown for discussions on LCBench runtime proxy, and Raul Astudillo for discussions on unknown cost handling for LogEIPC. AT thanks James T. Wilson for helpful comments on stopping rules for Bayesian optimization. LC is funded by the European Union (ERC-2022-SYG-OCEAN-101071601). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. AT was supported by Cornell University, jointly via the Center for Data Science for Enterprise and Society, the College of Engineering, and the Ann S. Bowers College of Computing and Information Science. PF was supported by AFOSR FA9550-20-1-0351. ZS was supported by the NSF under grant no. CMMI-2307008.
@@ Appendix B Theoretical Analysis and Calculations
In the following lemma, we prove a point-wise lower bound on the expected improvement before stopping for our recommended pairing of PBGI/LogEIPC stopping rule paired with either PBGI or LogEIPC acquisition function. [^4] In fact, any acquisition function (e.g., expected imporvement-cost (EIC) (Hu et al., 2025)) that ensures sufficient expected improvement relative to cost before the stopping time $\tau$ can apply here.
@@
PBGI. For each $x$ we defined a threshold $\alpha_{t}^{\mathrm{PBGI}}(x)$ by $\mathrm{EI}_{f\mid x_{1:t},y_{1:t}}\left(x;\alpha_{t}^{\mathrm{PBGI}}(x)\right)=c(x)$. Since $\mathrm{EI}_{f\mid x_{1:t},y_{1:t}}$ is increasing in its first argument,
@@
The existence of a point with ratio at least $1$ therefore implies that the set $\mathcal{S}_{t}=\{x:y_{1:t-1}^{*}\geq\alpha_{t}^{\mathrm{PBGI}}(x)\}$ is non-empty. PBGI chooses $x_{t}$ with the smallest threshold, and thus
$\alpha_{t}^{\mathrm{EI}}(x_{t})=\mathrm{EI}_{f\mid x_{1:t},y_{1:t}}\left(x;y_{1:t}^{*}\right)\geq\mathrm{EI}_{f\mid x_{1:t},y_{1:t}}\left(x;\alpha_{t}^{\mathrm{PBGI}}(x)\right)=c(x_{t}).$ (13)
(Log)EIPC. By definition $x_{t}$ maximizes $\log\left(\alpha_{t}^{\mathrm{EI}}(x)/c(x)\right)$, and thus $\alpha_{t}^{\mathrm{EI}}(x_{t})\geq c(x_{t})$.
@@
$\alpha_{t}^{\mathrm{EI}}(x_{t})\geq c(x_{t}),\qquad\text{for all }t\leq\tau.$ (14)
@@
We treat the two algorithms—meaning, the two acquisition function and stopping rule pairs—together and write $\mathcal{F}_{t}=\sigma\!\left(\{x_{i},y_{i}\}_{i=1}^{t}\right)$ for the filtration generated by the observations. Since we are minimizing, the one–step improvement after iteration $t$ is $y_{1:t-1}^{*}-y_{1:t}^{*}$, where recall that $y_{1:t}^{*}=\min_{1\leq i\leq t}y_{i}$. Since we initialize at the prior mean, we have $\mathbb{E}[y_{1:1}^{*}]=\mu$ and the quantity $y_{1:1}^{*}-\min\nolimits_{x\in X}f(x)$ has finite expectation $U<\infty$. Denote the posterior expected improvement function as $\alpha_{t}^{\mathrm{EI}}(x)=\mathrm{EI}_{f\mid x_{1:t},y_{1:t}}\left(x;y^{*}_{1:t}\right)$.
By Lemma˜1, $c(x_{t})\leq\alpha_{t}^{\mathrm{EI}}(x_{t})$ for all $t\leq\tau$. Set $\Delta_{t}=y_{1:t-1}^{*}-y_{1:t}^{*}\geq 0$ for $2\leq t\leq\tau$. Conditional on $\mathcal{F}_{t-1}$ and the choice of $x_{t}$, we have
$\mathbb{E}[\Delta_{t}\mid\mathcal{F}_{t-1},x_{t}]=\alpha_{t}^{\mathrm{EI}}(x_{t}).$ (15)
Taking expectations and summing up from 2 to $\tau$, by optional stopping theorem we get
$\mathbb{E}\left[\sum_{t=2}^{\tau}\alpha_{t}^{\mathrm{EI}}(x_{t})\right]=\mathbb{E}\left[\sum_{t=2}^{\tau}\Delta_{t}\right]=\mathbb{E}\left[y_{1:1}^{*}-y_{1:\tau}^{*}\right].$ (16)
Since we always have $y_{1:\tau}^{*}\geq\min_{x\in X}f(x)$, this gives
$\mathbb{E}\left[y_{1:1}^{*}-y_{1:\tau}^{*}\right]\leq\mu-\mathbb{E}\left[\min_{x\in X}f(x)\right]=U.$ (17)
@@
$\mathbb{E}\left[\sum_{t=2}^{\tau}c(x_{t})\right]\leq\mathbb{E}\left[\sum_{t=2}^{\tau}\alpha_{t}^{\mathrm{EI}}(x_{t})\right]\leq\mu-\mathbb{E}\left[\min_{x\in X}f(x)\right].$ (18)
We conclude that
$\mathbb{E}\left[\sum_{t=1}^{\tau}c(x_{t})\right]=\mathbb{E}[c(x_{1})]+\mathbb{E}\left[\sum_{t=2}^{\tau}c(x_{t})\right]\leq C+\mu-\mathbb{E}\left[\min_{x\in X}f(x)\right]=C+U,$ (19)
which completes the proof. ∎
Note 4.
In this paper, we focus on the setting where $f$ is drawn from a Gaussian process, i.e., $f\sim\mathcal{GP}(\mu,K)$. In this case, the term $U$ can be further bounded above using classical results on the expected supremum/infimum of Gaussian processes; see, for example, Lifshits (2012, Theorem 10.1).
@@
Since the Bayesian optimization algorithm considers the post-scaling cost, by Theorem˜2, $\mathbb{E}[\sum_{t=1}^{\tau}\lambda\cdot c(x_{t})]\leq\lambda C+U$. Since $\lambda=U/(B-C)$, we have
$\mathbb{E}\left[\sum_{t=1}^{\tau}c(x_{t})\right]\leq\frac{\lambda C+U}{\lambda}=C+\frac{U}{\lambda}=B.$

### v3

@@ 1 Introduction
In this work, we study how to design cost-aware stopping rules, motivated by two primary factors. First, state-of-the-art cost-aware acquisition functions such as the Pandora’s Box Gittins Index (PBGI) (Xie et al., 2024) and log expected improvement per cost (LogEIPC) (Ament et al., 2023) have not yet been studied in the adaptive stopping setting. This is important because—as our experiments in Section˜4 will show—for best performance, one should pair different acquisition functions with different stopping rules. Second, while certain stopping rules, such as UCB–LCB (Makarova et al., 2022), are guaranteed to achieve a low simple regret, they are not necessarily guaranteed to do so with low evaluation costs. This is important because—as our experiments will show—UCB–LCB will often incur high evaluation costs, resulting in a high cost-adjusted simple regret.
@@ 3.1 Theoretical Guarantees on Cost-Adjusted Simple Regret
Let $X$ be compact, and let $f:X\to\mathbb{R}$ be a random function with constant mean $\mu$. Consider a Bayesian optimization algorithm that begins at some initial point $x_{1}\in X$ with cost $C=c(x_{1})$, acquires additional points using either the PBGI or LogEIPC acquisition function, and uses the PBGI/LogEIPC stopping rule. Let $\tau=\min_{t\geq 1}\{\sup_{x\in X}\alpha_{t}^{\mathrm{LogEIPC}}(x)<0\}$ be the algorithm’s stopping time, and denote the posterior expected improvement function by $\alpha_{t}^{\mathrm{EI}}(x)=\mathrm{EI}_{f\mid x_{1:t},y_{1:t}}\left(x;y^{*}_{1:t}\right)$. Then, $\alpha_{t}^{\mathrm{EI}}(x_{t+1})\geq c(x_{t+1})$ for all $t<\tau$.
Proofs are in Appendix˜B. As a consequence of this claim, we show in Theorem˜2 that the PBGI/LogEIPC stopping rule, when paired with PBGI or LogEIPC, achieve strictly better cost-adjusted simple regret than the naive strategy of stopping immediately after the initial evaluation. While this may sound obvious, many acquisition function–stopping rule pairs fail to satisfy this no-worse-than-immediate property in practice (see Figures˜2, 3 and 4), largely because most stopping rules are designed with simple regret alone in mind and do not explicitly account for evaluation costs. To our knowledge, this is the first result establishing a formal guarantee on cost-adjusted simple regret for Bayesian optimization, generalizing beyond simpler settings such as Pandora’s Box, which assumes independent and discrete evaluations.
@@
Consider the setting, acquisition function, and stopping rule specified in Lemma˜1. Let $U=\mu-\mathbb{E}[\min_{x\in X}f(x)]<\infty$, then the algorithm’s expected cost-adjusted simple regret is bounded by
$\mathbb{E}\left[y_{1:\tau}^{*}-\min_{x\in X}f(x)+\sum_{t=1}^{\tau}c(x_{t})\right]\leq\mathbb{E}\Big[\underbrace{y_{1}-\min_{x\in X}f(x)+c(x_{1})}_{\textnormal{cost-adjusted regret of immediate stopping}}\Big]=U+C.$ (11)
@@ 4 Experiments
To evaluate our proposed PBGI/LogEIPC stopping rule, we design three complementary sets of experiments that progressively test its performance. First, we consider an idealized, low-dimensional Bayesian regret setting in which the Gaussian process model exactly matches the true objective. This controlled environment allows us to isolate the effect of different stopping rules without interference from modeling or optimization errors. Then, we move to a higher-dimensional Bayesian regret setting where each acquisition-function minimization must be approximated via a numerical optimizer. Finally, we test in practical scenarios with potential objective model mismatch, using the LCBench hyperparameter tuning benchmark and the NATS neural architecture size search benchmark. In each case, we compare pairs of acquisition functions and stopping rules, which we describe next.
@@
We compare the proposed PBGI/LogEIPC stopping rule against several stopping rules from prior work: UCB–LCB Makarova et al. (2022), LogEIPC-med Ishibashi et al. (2023), SRGap-med Ishibashi et al. (2023), and PRB Wilson (2024). We also include two simple heuristics used in practice: Convergence and GSS. UCB–LCB stops once the gap between upper and lower confidence bounds falls below a configurable threshold $\theta$. LogEIPC-med stops when the log expected improvement per cost drops beneath $\log(\eta)$ plus the median of its initial $I$ values. SRGap-med stops when the gap of the expected minimum simple regret falls below $\chi$ times the median of its initial $I$ values. PRB triggers once a probabilistic regret bound satisfies the regret tolerance $\epsilon$ and confidence $\delta$ parameters. Convergence stops as soon as the best observed value remains unchanged for $w$ iterations. GSS stops if the recent improvement is less than $\phi\times\mathrm{IQR}$ over the past $w$ trials where $\mathrm{IQR}$ denotes the inter-quartile range of current observations. Finally, we include two reference baselines: Immediate, which stops right after the initial evaluation (see Section˜3.1), and Hindsight, which, for each trial, selects the stopping time that yields the lowest cost-adjusted simple regret in hindsight, thus providing a lower bound on achievable performance.
@@ 4.1 Bayesian Regret
Figure 2: Cost-adjusted simple regret across acquisition-stopping rule pairs in 1D Bayesian regret setting. Objectives are sampled from a GP with Matérn-5/2 kernel, with a linear cost function scaled by $\lambda=0.1,0.01,0.001$. Matched PBGI pairs perform the best at $\lambda=0.1,0.01$, and remain close to hindsight-optimal at $\lambda=0.001$, though slightly outperformed by PRB and GSS paired with PBGI. The Immediate baseline is omitted at $\lambda=0.001$ due to its significantly higher regret (mean: 0.6942, error bar: [0.5314, 0.8570]) compared to other stopping rules.
Figure 3: Cost-adjusted simple regret across acquisition–stopping rule pairs in 8D Bayesian regret experiments. Objectives are sampled from a GP with a Matérn-5/2 kernel, using three different cost functions scaled by $\lambda=0.01$. Our PBGI/LogEIPC stopping rule consistently yields the lowest cost-adjusted regret that is close to hindsight optimal, when paired with LogEIPC, PBGI, and LCB.
@@ 4.2 Empirical Automated Machine Learning Benchmarks
Figure˜4 presents cost-adjusted regret for pairs of acquisition function and stopping rule on classification tasks from LCBench and NATS-Bench, with representative [^3] These were chosen to avoid degenerate cases—neither so large that the policy stops after only a few evaluations (e.g., $\lambda=10^{-4}$ on NATS-Bench, Figure 18) nor so small that evaluations are effectively free. cost-scaling factor $\lambda=10^{-4}$ and $\lambda=10^{-5}$ respectively. Additional results under different cost-scaling factors and the unknown-cost setting are provided in Figures˜15, 16 and 17 of Appendix˜D, along with visual comparisons between adaptive and fixed-iteration stopping. Overall, our PBGI/LogEIPC stopping rule consistently performs strongly in terms of cost-adjusted simple regret, particularly when paired with the PBGI acquisition function. We also report, in Table˜1 of Appendix˜D, how often each stopping rule fails to trigger within the 200-iteration cap: baselines such as SRGap-med and UCB–LCB often fail to stop early—particularly on NATS-Bench—whereas our stopping rule reliably stops before the cap.
@@
Figure 4: Comparison of cost-adjusted simple regret across acquisition function and stopping rule pairs on LCBench and NATS-Bench. The objective is to minimize validation error on classification tasks, with proxy runtime as evaluation cost, scaled by representative values of $\lambda$ ($10^{-4}$ for LCBench and $10^{-5}$ for NATS-Bench); see Figures˜12, 13 and 14 for performance under other $\lambda$. The matched PBGI combination performs best on adult, higgs, and cifar10-valid, second-best on Fashion-MNIST, and third-best on the remaining two, while closely approaching the hindsight optimal across all tasks. Our stopping rule is also competitive when paired with LogEIPC or TS.
@@ Acknowledgments
QX thanks Nairen Cao for helpful discussions on hindsight optimal, Theodore Brown for helpful discussions on LCBench runtime proxy, and Raul Astudillo for helpful discussions on unknown cost handling for LogEIPC. AT thanks James T. Wilson for helpful comments on stopping rules for Bayesian optimization. LC is funded by the European Union (ERC-2022-SYG-OCEAN-101071601). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. AT was supported by Cornell University, jointly via the Center for Data Science for Enterprise and Society, the College of Engineering, and the Ann S. Bowers College of Computing and Information Science. PF was supported by AFOSR FA9550-20-1-0351. ZS was supported by the NSF under grant no. CMMI-2307008.
@@ Appendix B Theoretical Analysis and Calculations
In the following lemma, we prove a point-wise lower bound on the expected improvement before stopping for our recommended pairing of PBGI/LogEIPC stopping rule paired with either PBGI or LogEIPC acquisition function. [^4] In fact, any acquisition function (e.g., expected improvement-cost (EIC) (Hu et al., 2025)) that ensures the one-step expected improvement is worth the evaluation cost before the stopping time $\tau$ can apply here.
@@
PBGI. For each $x$ we defined a threshold $\alpha_{t}^{\mathrm{PBGI}}(x)$ by $\mathrm{EI}_{f\mid x_{1:t},y_{1:t}}\left(x;\alpha_{t}^{\mathrm{PBGI}}(x)\right)=c(x)$. Since $\mathrm{EI}_{f\mid x_{1:t},y_{1:t}}$ is increasing in its second argument,
@@
The existence of a point with ratio at least $1$ therefore implies that the set $\mathcal{S}_{t}=\{x:y_{1:t}^{*}\geq\alpha_{t}^{\mathrm{PBGI}}(x)\}$ is non-empty. PBGI chooses $x_{t+1}$ with the smallest threshold, and thus
$\alpha_{t}^{\mathrm{EI}}(x_{t+1})=\mathrm{EI}_{f\mid x_{1:t},y_{1:t}}\left(x_{t+1};y_{1:t}^{*}\right)\geq\mathrm{EI}_{f\mid x_{1:t},y_{1:t}}\left(x_{t+1};\alpha_{t}^{\mathrm{PBGI}}(x_{t+1})\right)=c(x_{t+1}).$ (13)
(Log)EIPC. By definition $x_{t+1}$ maximizes $\log(\alpha_{t}^{\mathrm{EI}}(x)/c(x))$ and $\alpha_{t}^{\mathrm{EI}}(x)/c(x)$, hence $\alpha_{t}^{\mathrm{EI}}(x_{t+1})\geq c(x_{t+1})$.
@@
$\alpha_{t}^{\mathrm{EI}}(x_{t+1})\geq c(x_{t+1}),\qquad\text{for all }t<\tau.$ (14)
@@
We treat the two algorithms—meaning, the two acquisition function and stopping rule pairs—together and write $\mathcal{F}_{t}=\sigma\!\left(\{x_{i},y_{i}\}_{i=1}^{t}\right)$ for the filtration generated by the observations. Since we are minimizing, the one–step improvement after iteration $t$ is $y_{1:t-1}^{*}-y_{1:t}^{*}$, where recall that $y_{1:t}^{*}=\min_{1\leq i\leq t}y_{i}$. By our assumption about the random function $f$, $\mathbb{E}[y_{1:1}^{*}]=\mu$ and the quantity $y_{1:1}^{*}-\min\nolimits_{x\in X}f(x)$ has finite expectation $U<\infty$. Denote the posterior expected improvement function as $\alpha_{t}^{\mathrm{EI}}(x)=\mathrm{EI}_{f\mid x_{1:t},y_{1:t}}\left(x;y^{*}_{1:t}\right)$.
By Lemma˜1, $c(x_{t})\leq\alpha_{t-1}^{\mathrm{EI}}(x_{t})$ for all $t\leq\tau$. Set $\Delta_{t}=y_{1:t-1}^{*}-y_{1:t}^{*}\geq 0$ for $2\leq t\leq\tau$. Conditional on $\mathcal{F}_{t-1}$ and the choice of $x_{t}$, we have
$\mathbb{E}[\Delta_{t}\mid\mathcal{F}_{t-1},x_{t}]=\alpha_{t-1}^{\mathrm{EI}}(x_{t}).$ (15)
Taking expectations and summing up from 2 to $\tau$,
$\mathbb{E}\!\left[\sum_{t=2}^{\tau}\alpha^{\mathrm{EI}}_{t-1}(x_{t})\right]$ $=\mathbb{E}\!\left[\sum_{t=2}^{\infty}\mathbf{1}_{\{t\leq\tau\}}\,\alpha^{\mathrm{EI}}_{t-1}(x_{t})\right]$
$=\mathbb{E}\!\left[\sum_{t=2}^{\infty}\mathbf{1}_{\{t\leq\tau\}}\,\mathbb{E}\!\left[\Delta_{t}\mid\mathcal{F}_{t-1},x_{t}\right]\right]$
$=\mathbb{E}\!\left[\sum_{t=2}^{\infty}\mathbf{1}_{\{t\leq\tau\}}\,\Delta_{t}\right]\quad(\text{tower property})$
$=\mathbb{E}\!\left[\sum_{t=2}^{\tau}\Delta_{t}\right]=\mathbb{E}\!\left[y^{*}_{1:1}-y^{*}_{1:\tau}\right].$ (16)
@@
$\mathbb{E}\left[\sum_{t=2}^{\tau}c(x_{t})\right]\leq\mathbb{E}\left[\sum_{t=2}^{\tau}\alpha_{t-1}^{\mathrm{EI}}(x_{t})\right]\leq\mathbb{E}\!\left[y^{*}_{1:1}-y^{*}_{1:\tau}\right].$ (17)
Adding the term $y^{*}_{1:\tau}$ on both sides of (17) gives
$\mathbb{E}\!\left[y^{*}_{1:\tau}+\sum_{t=2}^{\tau}c(x_{t})\right]\leq\mathbb{E}[\,y^{*}_{1:1}\,].$
Finally, adding $c(x_{1})$ to both sides and subtracting $\mathbb{E}\!\left[\min_{x\in X}f(x)\right]$ yields
$\mathbb{E}\!\left[y^{*}_{1:\tau}-\min_{x\in X}f(x)+\sum_{t=1}^{\tau}c(x_{t})\right]$ $\leq\mathbb{E}\!\left[y^{*}_{1:1}-\min_{x\in X}f(x)+c(x_{1})\right]$
$=\mathbb{E}\!\left[y_{1}-\min_{x\in X}f(x)+c(x_{1})\right]$
$=\mu-\mathbb{E}\!\left[\min_{x\in X}f(x)\right]+\mathbb{E}[c(x_{1})]$
$=U+C,$
since $\mathbb{E}[y_{1}]=\mu$, $C=c(x_{1})$, and $U=\mu-\mathbb{E}\left[\min_{x\in X}f(x)\right]$. This is exactly (11). ∎
@@
Since the Bayesian optimization algorithm considers the post-scaling cost, by Theorem˜2 and the fact that $y^{*}_{1:\tau}-\min_{x\in X}f(x)>0$ pointwise, we have
$\mathbb{E}\left[\sum_{t=1}^{\tau}\lambda\cdot c(x_{t})\right]\leq\mathbb{E}\!\left[y^{*}_{1:\tau}-\min_{x\in X}f(x)+\sum_{t=1}^{\tau}\lambda\cdot c(x_{t})\right]\leq\mu-\mathbb{E}\left[\min_{x\in X}f(x)\right]+\lambda C=U+\lambda C.$ (18)
Since $\lambda=U/(B-C)$, we have
$\mathbb{E}\left[\sum_{t=1}^{\tau}c(x_{t})\right]\leq\frac{U+\lambda C}{\lambda}=\frac{U}{\lambda}+C=B.$
Note 4.
In this paper, we focus on the setting where $f$ is drawn from a Gaussian process, i.e., $f\sim\mathcal{GP}(\mu,K)$. In this case, the term $U=\mu-\mathbb{E}\!\left[\min_{x\in X}f(x)\right]$ can be further bounded above using classical results on the expected supremum/infimum of Gaussian processes; see, for example, Lifshits (2012, Theorem 10.1).